OpenStudy (mathteacher1729):

THIS JUST HAPPENED -- ("real world math") Seven co-workers take one friend out to lunch on her last day of work before she moves to a different state. The bill comes to $96.94, BEFORE an 18% tip is applied. The friend who is moving does NOT pay, and her cost is distributed evenly among the others. When the bill arrives $85 is paid in cash plus two credit cards, one paying $15, the other $12. The waitress argues that there was not enough money paid. Is the waitress correct, and if so, how much was underpaid? Explain your reasoning to the angry waitress. ;-p

OpenStudy (anonymous):

i think enough is paid! how much does the waitress want in tips? XD

OpenStudy (mathteacher1729):

Why do you think enough was paid? (Was the exact amount paid or did the group overpay?) This is one of those cases where "Showing your work" in real life is quite important. :-p

OpenStudy (anonymous):

Well the total with tax would be 114.04 they paid 85+15+12=112

OpenStudy (anonymous):

So $2.04 short i think?

OpenStudy (anonymous):

ooooooooh! i see. so the waiter did have the right :)

OpenStudy (mathteacher1729):

@CalebBeavers -- correct numbers. For those who may not understand where $114.04 came from, how did you obtain that number?

OpenStudy (anonymous):

because you added everything that they paid... ?

OpenStudy (anonymous):

18% of 96.94 added ontop of 96.94

OpenStudy (anonymous):

.18*96.64+96.64

OpenStudy (anonymous):

96.64 x .18 = 17.45 (18% tip) 96.64 + 17.45 = 114.04

OpenStudy (mathteacher1729):

@finkimlost and @CalebBeavers thank you for clarifying. :) You can also do: 96.64*1.18 = 114.04

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

that too lol